# Adver-City: Open-Source Multi-Modal Dataset for Collaborative Perception Under Adverse Weather Conditions

### v1

@@ Abstract
Adverse weather conditions pose a significant challenge to the widespread adoption of Autonomous Vehicles (AVs) by impacting sensors like LiDARs and cameras. Even though Collaborative Perception (CP) improves AV perception in difficult conditions, existing CP datasets lack adverse weather conditions. To address this, we introduce Adver-City, the first open-source synthetic CP dataset focused on adverse weather conditions. Simulated in CARLA with OpenCDA, it contains over $24$ thousand frames, over $890$ thousand annotations, and $110$ unique scenarios across six different weather conditions: clear weather, soft rain, heavy rain, fog, foggy heavy rain and, for the first time in a synthetic CP dataset, glare. It has six object categories including pedestrians and cyclists, and uses data from vehicles and roadside units featuring LiDARs, RGB and semantic segmentation cameras, GNSS, and IMUs. Its scenarios, based on real crash reports, depict the most relevant road configurations for adverse weather and poor visibility conditions, varying in object density, with both dense and sparse scenes, allowing for novel testing conditions of CP models. Benchmarks run on the dataset show that weather conditions created challenging conditions for perception models, reducing multi-modal object detection performance by up to $19\%$, while object density affected LiDAR-based detection by up to $29\%$. The dataset, code and documentation are available at https://labs.cs.queensu.ca/quarrg/datasets/adver-city/.
@@
I Introduction
Self-driving cars are here, with robotaxi companies in the USA and China having already logged over $60$ million fully autonomous miles [1]. However, Autonomous Vehicles (AVs) still have a long road ahead before reaching widespread adoption. Many issues must still be addressed to ensure AVs can drive safely and reliably in all possible scenarios.
Adverse weather conditions such as rain, fog, and glare present significant challenges for AVs. Although these conditions do not constitute the most common driving conditions, they create particularly difficult driving scenarios. For instance, it only rains over land $8\%$ of the time [2] but the risk of accidents in rainy weather is $70\%$ higher than normal [3]. Additionally, these conditions degrade the performance of perception sensors like cameras and LiDARs [4].
However, the field faces a shortage of adverse weather datasets, as there is currently no open-source CP dataset focused on such conditions. Existing open-source datasets have no adverse weather scenes, while the ones that have adverse conditions offer little weather diversity and are not open-source, preventing scene modifications to expand weather variety.
We introduce Adver-City, the first open-source CP dataset focused on adverse weather conditions, created using the CARLA simulator [8] and OpenCDA framework [9]. The dataset includes $110$ diverse scenarios across six distinct weather conditions: clear weather, glare, soft rain, heavy rain, fog and foggy heavy rain (Fig. 1). It is the first synthetic CP dataset to feature glare and incorporate real-life data on adverse weather and poor visibility conditions in its scenario design. Each scene includes three CAVs and two RSUs, and multi-modal data from LiDARs, RGB and semantic cameras, GNSS, and IMUs, supporting tasks like object detection and tracking, as well as semantic segmentation.
@@ I Introduction
This paper is organized as follows: Section II presents and discusses synthetic multi-modal CP datasets, Section III describes the Adver-City dataset in regards to its scenarios, generation and statistics, Section IV presents the results of benchmarks performed with the dataset, and Section V discusses possible directions to be taken in the future.
II Related Work
Collaborative Perception has gained considerable attention recently, leading to the development of various datasets. Some, like DAIR-V2X [10], V2V4Real [11] and TUMTraf [12], are based on real-world data, while others use synthetic data from simulators such as CARLA [8] and LidarSIM [13]. Among the synthetic datasets, some focus solely on LiDAR data, such as V2V-Sim [14] and CODD [15], while others provide multi-modal data from both LiDARs and cameras, as shown in Table I and discussed here.
@@ II Related Work
TABLE I: Comparison between multi-modal CP datasets generated using CARLA [8]. The number of annotations for V2X-Sim and DeepAccident is not provided by their authors. All datasets are publicly available, yet only the Open Source datasets provide scenario generation code.
OPV2V [16] is arguably the most used CP dataset, serving as a benchmark for several collaborative models, including Coalign [20], Where2Comm [21] and CoBEVT [22]. Built with the open-source OpenCDA framework [9], it has inspired variations such as OPV2V-H [23], OPV2V+ [24] and OPV2VH+ [25]. However, it does not contain RSUs, includes only one object category (cars), and features only clear weather scenarios.
@@
The DOLPHINS [18] dataset, while open-source, is less commonly used and built on a less robust codebase than OpenCDA’s [9]. Additionally, it includes only six clear weather scenarios with data captured from a LiDAR sensor and a single front-facing camera per agent.
Focusing on accident prediction, DeepAccident [7] features a large number of scenarios with diverse object categories, including pedestrians and two-wheeled vehicles. The scenes, selected based on real-world crash reports, include a few environmental conditions – $37\%$ are rainy, and $16.3\%$ occur at night. However, the scenario generation code is not publicly available, and, due to its focus on accident prediction, the scenes are short, averaging less than $9$ seconds in length.
The recent SCOPE dataset [19] performs image and point cloud augmentation on top of CARLA’s sensor data to provide more realistic weather. It includes data from digital twin maps of German cities and a solid-state LiDAR among its sensors. Despite its focus on adverse weather, only $44.2\%$ of its scenes feature fog or rain, and just $25.8\%$ take place at night. Additionally, the scenario generation code has not been made publicly available.
III Adver-City Dataset
III-A Scenarios
Inspired by DeepAccident [7], we selected scenarios based on comprehensive real-world crash data from the US National Highway Traffic Safety Administration (NHTSA). Even though accidents involving AVs are fundamentally different from those involving only non-autonomous vehicles (over $90\%$ of crashes are caused by driver error [30]), data on AV crashes remains scarce. Public AV crash reports document only $30$ [31] or fewer [32] AV crashes, compared to the latest NHTSA report [33], which records over $5$ million accidents for non-autonomous vehicles.
An NHTSA report from 2007 [34] defined a typology consisting of $37$ pre-crash scenarios based on vehicle movements and the critical event occurring immediately before a crash, presenting detailed statistics on each scenario. However, in 2019, another NHTSA report [33] updated the typology to a total of $36$ pre-crash scenarios, clustering them in $9$ scenario groups (Table II) and providing up-to-date statistics on each group.
@@ III-A Scenarios
TABLE II: Statistics from the latest NHTSA report [33] on pre-crash scenario groups. The highest two values in each column are highlighted for convenience. *Percentage of crashes in each scenario group that had adverse conditions as the main factor. Average number of vehicles per crash in each group.
Since the main problems for object detection under adverse weather conditions may be attributed to poor visibility and sensor performance degradation [35], we filtered the pre-crash scenarios groups by focusing on those that have, among their main causes, errors in perception (i.e., identification of an object, such as a car, a pedestrian or a cyclist) instead of errors in control (often caused by driver impairment, driver inexperience or speeding). Moreover, we focused on scenario groups with the highest fatality rates, strongly influenced by adverse weather conditions and poor visibility.
@@
In light of this data, we selected five road configurations that cover all of the above cases and modelled them on the following CARLA maps [8]:
@@
  • Light rain (Day and Night);
@@
By combining $11$ weather and $2$ density levels across $5$ road configurations, we created a total of $110$ unique scenarios. This setup enables detailed comparisons between environmental factors, such as the impact of rain intensity across different locations or the effects of increased object density at intersections. Despite stemming from only five road configurations, each scenario was simulated independently so that, for a given road configuration and density pairing, the agents’ behaviour varies slightly, resulting in a diverse set of simulations.
III-B Dataset Generation
The Adver-City dataset was generated in the CARLA simulator [8] with the OpenCDA [9] framework. Each scenario has data from five distinct viewpoints, as illustrated in Fig. 2: two RSUs, the ego vehicle, and two CAVs. Each viewpoint is equipped with the sensors listed in Table III.
@@ III-B Dataset Generation
TABLE III: Sensor specifications for the AVs used in Adver-City scenarios. The RSUs are not equipped with GNSS & IMU but otherwise follow the same specification. *RSUs positioned above ground have a $-25^{\circ}$ to $-1^{\circ}$ vertical FOV.
The sensor configuration was selected to balance detection performance with real-world feasibility. We used $1080$p ($1920\times 1080$) cameras and $32$-channel LiDARs while omitting depth cameras. The $1080$p cameras offer more detailed imagery compared to lower resolution options and are relatively affordable and widely available. The $32$-channel LiDARs offer less detail than $64$-channel models but are also more accessible. Although depth cameras may provide unique information, they are rarely used in practice. None of the real-world datasets discussed in Section II use depth cameras; instead, they use LiDARs with fewer than $64$ channels and cameras with a resolution of $1080$p or higher.
The AVs’ sensors are positioned as illustrated in Fig. 3: the frontal camera is mounted on the windshield, the rear camera is located at the back of the trunk, side-facing cameras are placed at the top of the B-pillars (tilted backwards by $10^{\circ}$ to reduce motion blur), and the LiDAR is positioned at the top of the vehicle. RSUs are either attached to tall structures (such as walls, towers, traffic light poles, or light posts) or positioned in high-visibility spots along the road. In non-junction scenarios, one CAV is positioned ahead of the ego vehicle, travelling in the same direction, while the other CAV is also ahead but travelling in the opposite direction. In intersection scenarios, CAVs approach the intersection from different roads compared to the ego vehicle.
@@
Data is captured at $10$Hz and adheres to an annotation schema compatible with OPV2V [16], with the main difference relating to the categories in our dataset. Unlike OPV2V, which only included cars and did not require category annotations for bounding boxes, our dataset has six categories: cars, pedestrians, vans, trucks, bicycles, and motorcycles.
III-C Dataset Statistics
Fig. 4 presents statistics for the Adver-City dataset. On average, scenarios last $21.9$ seconds ($219$ frames at $10$Hz) and contain $52$ unique vehicles and $44$ unique pedestrians. Overall, the dataset has $24,087$ frames and $890,127$ annotations pertaining to $11,418$ vehicles and $4,840$ pedestrians.
@@ III-C Dataset Statistics
Fig. 4(a) and Fig. 4(b) highlight the diversity of environmental conditions on Adver-City, showing slight variations in the percentage of frames for each weather condition. These differences result from running scenarios individually, leading to minor variations in agent behaviour and consequently in scenario length. Fig. 4(c) shows the impact of density settings, with dense scenarios having significantly more objects, especially near the ego vehicle, compared to sparse ones.
@@
Fig. 4(d) and Fig. 4(e) showcase the diversity of frames in Adver-City. The first shows the balance between the number of annotations in the keyframes, ranging from just a few to over $120$ bounding boxes. The second shows that objects are positioned at various angles and distances from the ego vehicle, with a concentration around axial angles due to road layout. Fig. 4(f) highlights that the Pedestrian category has the highest number of annotated instances, making Adver-City well-suited for detecting Vulnerable Road Users (VRUs).
IV Benchmark
@@
How can a collaborative 3D object detection model trained without adverse weather generalize to adverse weather conditions? To answer that question, we used models trained on OPV2V [16] (which has no adverse weather) to perform inference on Adver-City, having as a baseline the Culver City testing scenario from OPV2V.
We tested the trained baseline models provided by the authors of HEAL [23] on Culver City, selecting the best-performing one for each sensor modality. Since the models for camera-based detection were trained with depth data, they were disregarded from our analysis as the removal of depth made the models unsuitable for detection (the top camera-based model achieved an AP@$50$ of $19.02$). Both for LiDAR and LiDAR + Camera detection, CoBEVT [22] was the best model for Culver City and was then used for inference in the Adver-City scenarios presented in Table IV. PointPillars [38] was used as the backbone for LiDAR models, and Lift-Splat-Shoot [39], with EfficientNet [40] as the image encoder, was used for camera experiments. All Adver-City experiments were conducted across the five road configurations described in Section III-A.
@@
Even though our dataset follows the OPV2V annotation schema, minor adjustments were required to ensure the OPV2V-trained model performed adequately on Adver-City. Since OPV2V includes only the ‘car’ category, the model was limited to detecting cars. Bounding boxes for dissimilar categories such as ‘pedestrian’, ‘bicycle’, and ‘motorcycle’, were removed for the benchmark. Also, data from RSUs was excluded since OPV2V does not support V2X collaboration.
Dataset Scenario Lidar (AP@50) Lidar + Camera (AP@50)

### v2

@@ Abstract
Adverse weather conditions pose a significant challenge to the widespread adoption of Autonomous Vehicles (AVs) by impacting sensors like LiDARs and cameras. Even though Collaborative Perception (CP) improves AV perception in difficult conditions, existing CP datasets lack adverse weather conditions. To address this, we introduce Adver-City, the first open-source synthetic CP dataset focused on adverse weather conditions. Simulated in CARLA with OpenCDA, it contains over $24$ thousand frames, over $890$ thousand annotations, and $110$ unique scenarios across six different weather conditions: clear weather, soft rain, heavy rain, fog, foggy heavy rain and, for the first time in a synthetic CP dataset, glare. It has six object categories including pedestrians and cyclists, and uses data from vehicles and roadside units featuring LiDARs, RGB and semantic segmentation cameras, GNSS, and IMUs. Its scenarios, based on real crash reports, depict the most relevant road configurations for adverse weather and poor visibility conditions, varying in object density, with both dense and sparse scenes, allowing for novel testing conditions of CP models. Benchmarks run on the dataset show that weather conditions created challenging conditions for perception models, with CoBEVT scoring $58.30/52.44/38.90$ (AP@30/50/70). The dataset, code and documentation are available at https://labs.cs.queensu.ca/quarrg/datasets/adver-city/.
1 Introduction
Self-driving cars are here, with robotaxi companies in the USA and China having already logged over $60$ million fully autonomous miles on public roads [3]. However, Autonomous Vehicles (AVs) still have a long road ahead before reaching widespread adoption. Many issues must still be addressed to ensure that AVs can drive safely and reliably in all possible scenarios.
Adverse weather conditions such as rain, fog, and glare present significant challenges to AVs. Although most driving is not done in these conditions, they create particularly difficult driving scenarios. For instance, it only rains on land $8\%$ of the time [36] but the risk of accidents in rainy weather is $70\%$ higher than normal [1]. Furthermore, adverse weather degrades the performance of perception sensors such as cameras and LiDARs [26].
However, the development of robust CP models for challenging weather is hindered by the fact that no public real-world CP dataset focuses on adverse weather conditions. Despite their reduced realism, simulated datasets offer a promising solution to such a shortage through controlled scenario design, allowing uncommon scenarios such as heavy fog and intense glare to be captured extensively. Yet, existing open-source CP datasets have no adverse weather scenes, while those with adverse conditions offer little weather diversity and are not open-source, preventing scene modifications to expand weather variety.
@@ 1 Introduction
To address this gap, we introduce Adver-City, the first open-source CP dataset focused on adverse weather, created using the CARLA simulator [6] and OpenCDA framework [40]. The dataset includes $110$ diverse scenarios across six distinct weather conditions: clear weather, glare, soft rain, heavy rain, fog and foggy heavy rain (Figure 1). It is the first synthetic CP dataset to feature glare and incorporate real-life data on adverse weather and poor visibility conditions in its scenario design. Each scene includes three CAVs and two RSUs, and multi-modal data from LiDARs, RGB and semantic cameras, GNSS, and IMUs, supporting tasks like object detection, tracking, and semantic segmentation.
@@
This paper is organized as follows: Section 2 presents and discusses synthetic multi-modal CP datasets, Section 3 describes the Adver-City dataset in regard to its scenarios, generation and statistics, Section 4 presents the results of benchmarks performed with the dataset, Section 5 discusses limitations regarding the realism of the dataset, and Section 6 discusses possible directions to be taken in the future.
2 Related Work
Collaborative Perception has recently gained considerable attention, leading to the development of various datasets. Some, like DAIR-V2X [45], V2V4Real [43] and TUMTraf [47], are based on real-world data, while others use synthetic data from simulators such as CARLA [6] and LidarSIM [23]. Among the synthetic datasets, some focus solely on LiDAR data, such as V2V-Sim [39] and CODD [2], while others provide multi-modal data from both LiDARs and cameras, as shown in Table 1 and discussed here.
@@ 2 Related Work
Table 1: Comparison between multi-modal CP datasets generated using CARLA [6]. The number of annotations for V2X-Sim and DeepAccident is not provided by their authors. All datasets are publicly available, yet only the Open Source datasets provide scenario generation code.
OPV2V [41] is arguably the most used CP dataset to date, serving as a benchmark for several collaborative models, including Coalign [20], Where2Comm [10] and CoBEVT [42]. Built with the open-source OpenCDA framework [40], it has inspired variations such as OPV2V-H [21], OPV2V+ [11] and OPV2VH+ [12]. However, it does not contain RSUs, includes only one object category (cars), and features only clear weather scenarios.
@@
The DOLPHINS [24] dataset, while open-source, is less commonly used and built on a less robust codebase than OpenCDA’s [40]. In addition, it includes only six clear weather scenarios with data captured from a LiDAR sensor and a single front-facing camera per agent.
Focusing on accident prediction, DeepAccident [38] features a large number of scenarios with diverse object categories, including pedestrians and two-wheeled vehicles. The scenes, selected based on real-world crash reports, include a few environmental conditions – $37\%$ are rainy and $16.3\%$ occur at night. However, the scenario generation code is not publicly available and, due to its focus on accident prediction, the scenes are short, averaging less than $9$ seconds in length.
The recent SCOPE dataset [7] performs image and point cloud augmentation on top of CARLA’s sensor data to provide more realistic weather. It includes data from digital twin maps of German cities and a solid-state LiDAR among its sensors. Despite its focus on adverse weather, only $44.2\%$ of its scenes feature fog or rain, and only $25.8\%$ take place at night. Additionally, neither the scenario generation code nor the weather augmentation algorithms have been made publicly available.
3 Adver-City Dataset
3.1 Scenarios
Inspired by DeepAccident [38], we selected scenarios based on comprehensive real-world crash data from the US National Highway Traffic Safety Administration (NHTSA). Even though accidents involving AVs are fundamentally different from those involving only non-autonomous vehicles (over $90\%$ of crashes are caused by driver error [32]), data on AV crashes remain scarce. Public AV crash reports document only $30$ [19] or fewer [18] AV crashes, compared to the latest NHTSA report [33], which records over $5$ million accidents for non-autonomous vehicles.
An NHTSA report from 2007 [25] defined a typology consisting of $37$ pre-crash scenarios based on vehicle movements and the critical event occurring immediately before a crash, presenting detailed statistics on each scenario. Then, in 2019, another NHTSA report [33] updated the typology to a total of $36$ pre-crash scenarios, clustering them in $9$ scenario groups (Table 2) and providing up-to-date statistics on each group.
@@ 3.1 Scenarios
Table 2: Statistics from the latest NHTSA report [33] on pre-crash scenario groups. The highest two values in each column are highlighted for convenience. *Percentage of crashes in each scenario group that had adverse conditions as the main factor. Average number of vehicles per crash in each group.
Since the main problems for object detection under adverse weather conditions may be attributed to poor visibility and sensor performance degradation [30], we filtered the pre-crash scenarios groups by focusing on those that have, among their main causes, errors in perception (i.e. identification of an object, such as a car or pedestrian) instead of errors in control (often caused by driver impairment, driver inexperience or speeding). Moreover, we focused on scenario groups with the highest fatality rates, strongly influenced by adverse weather conditions and poor visibility.
@@
In light of this data, we selected five road configurations that cover all of the above cases and modeled them on the following CARLA maps [6]:
@@
  • Soft rain (Day and Night);
@@
By combining $11$ weather and $2$ density levels across $5$ road configurations, we created a total of $110$ unique scenarios. This setup enables detailed comparisons between environmental factors, such as the impact of rain intensity across different locations or the effects of increased object density at intersections. Despite stemming from only five road configurations, each scenario was simulated independently so that, for a given road configuration and density pairing, the agents’ behavior varies slightly, resulting in a diverse set of simulations.
3.2 Dataset Generation
The Adver-City dataset was generated in the CARLA simulator [6] with the OpenCDA [40] framework. Each scenario has data from five distinct viewpoints, as illustrated in Figure 2: two RSUs, the ego vehicle, and two CAVs. Each viewpoint is equipped with the sensors listed in Table 3.
@@ 3.2 Dataset Generation
Table 3: Sensor specifications for the AVs used in Adver-City scenarios. The RSUs are not equipped with GNSS & IMU but otherwise follow the same specification. *RSUs positioned above ground have a $-25^{\circ}$ to $-1^{\circ}$ vertical FOV.
The sensor configuration was selected to balance detection performance with real-world feasibility. We used $1080$p ($1920\times 1080$) cameras and $32$-channel LiDARs while omitting depth cameras. The $1080$p cameras offer more detailed imagery compared to lower resolution options and are relatively affordable and widely available. The $32$-channel LiDARs offer less detail than $64$-channel models but are also more accessible. Although depth cameras may provide unique information, they are rarely used in practice. None of the real-world datasets discussed in Section 2 use depth cameras; instead, they use LiDARs with fewer than $64$ channels and cameras with a resolution of $1080$p or higher.
The AVs’ sensors are positioned as illustrated in Figure 3: the frontal camera is mounted on the windshield, the rear camera is located at the back of the trunk, side-facing cameras are placed at the top of the B-pillars (tilted backwards by $10^{\circ}$ to reduce motion blur), and the LiDAR is positioned at the top of the vehicle. RSUs are either attached to tall structures (such as walls, towers, traffic light poles, or light posts) or positioned in high-visibility spots along the road. In non-junction scenarios, one CAV is positioned ahead of the ego vehicle, travelling in the same direction, while the other CAV is also ahead but travelling in the opposite direction. In intersection scenarios, CAVs approach the intersection from different roads compared to the ego vehicle. More details on the placement of RSUs and CAVs for each road configuration are provided in the Supplementary Material.
@@
Data was captured at $10$Hz and adheres to an annotation schema compatible with OPV2V [41], with the main difference relating to the categories in our dataset. Unlike OPV2V, which only included cars and did not require category annotations for bounding boxes, our dataset has six categories: cars, pedestrians, vans, trucks, bicycles, and motorcycles.
3.3 Dataset Statistics
Figure 4 presents statistics for the Adver-City dataset. On average, scenarios last $21.9$ seconds ($219$ frames at $10$Hz) and contain $52$ unique vehicles and $44$ unique pedestrians. Overall, the dataset has $24,087$ frames and $890,127$ annotations pertaining to $11,418$ vehicles and $4,840$ pedestrians.
@@ 3.3 Dataset Statistics
Figure 4(a) and Figure 4(b) highlight the diversity of environmental conditions in Adver-City, showing slight variations in the percentage of frames for each weather condition. These differences result from running scenarios individually, leading to minor variations in agent behaviour and consequently in scenario length. Figure 4(c) shows the impact of density settings, with dense scenarios having significantly more objects, especially near the ego vehicle, compared to sparse ones.
@@
Figure 4(d) and Figure 4(e) showcase the diversity of frames in Adver-City. The first shows the balance between the number of annotations in the keyframes, ranging from just a few to over $120$ bounding boxes. The second shows that objects are positioned at various angles and distances from the ego vehicle, with a concentration around axial angles due to road layout. Figure 4(f) highlights that the Pedestrian category has the highest number of annotated instances, making Adver-City well-suited for detecting vulnerable road users.
4 Benchmark
@@
How can a collaborative multi-modal 3D object detection model trained without adverse weather generalize to adverse weather conditions? To answer that question, we used the CoBEVT [42] model trained on OPV2V [41] (which has no adverse weather) to perform inference on Adver-City, having as a baseline the Culver City test scenario from OPV2V.
We used the trained model provided by the authors of HEAL [21], which has PointPillars [13] as the backbone for LiDAR, and Lift-Splat-Shoot [28] for camera with EfficientNet [35] as the image encoder.
@@
Even though our dataset follows the OPV2V annotation schema, minor adjustments were required to ensure the OPV2V-trained model performed adequately on Adver-City. Since OPV2V includes only the ‘car’ category, the model was limited to detecting cars. Bounding boxes for dissimilar categories such as ‘pedestrian’, ‘bicycle’, and ‘motorcycle’, were removed from the experiment. Also, data from RSUs was excluded since OPV2V does not support V2X collaboration.
Then, to get a benchmark of the dataset, we took the same CoBEVT model (pre-trained on OPV2V) and trained it for 20 more epochs on Adver-City with a learning rate of $0.00001$. We split our dataset by road configurations, assigning the rural intersection, rural straight non-junction and urban non-junction scenarios to the training set, rural curved non-junction to the validation set, and urban intersection to the test set. As such, each set presents all weather conditions and density settings while avoiding data leakage. The resulting frame count for the train/validation/test split was $14,910/4,774/4,403$. Since further training was performed on the model, we included the data from the RSUs, but kept the ‘pedestrian’, ‘bicycle’, and ‘motorcycle’ categories out of the experiment to guarantee a fair comparison with the OPV2V-trained model.
The results of both experiments are presented in Table 4. For both, detection was evaluated for objects near the ego vehicle in a range of $x\in[-102.4,102.4]$, $y\in[-102.4,102.4]$, with CAVs having a communication range of $70$ meters. For the OPV2V-trained model, we performed inference on the entire Adver-City dataset, with the five road configurations described in Section 3.1 present in each combination of adverse weather conditions and density settings.
Dataset Scenario
Trained on OPV2V
(LiDAR + Camera, AP@50)
Trained on Adver-City
(LiDAR + Camera, AP@50)